OpenStudy (anonymous):

if the length of a pool is 50m , its width is 25m and height is 4m.we r required to paint its wall and floor what's the expenditures if the cost of paint per meter is 25dollars. solve it with proper steps.

OpenStudy (anonymous):

 Area of walls and floor is 2a + 2b + c. 2(a+b)+c = 8(25 + 50) + (50x25) = 1850m^2